I need your help.. thank you.
The Herbalist.
A certain farmer sold one-third of his crop to miss liberty, one fourth to freewheeling Franklin, one -sixth to Daffy Duck, one-tenth to Tony Soprano, and twelfth to J.R. 'Bob' Dobbs; he have one -twentieth to lady friend, and he consumed the last ounce himself. How much did he grow, and how much did each person get?

Answer by richwmiller(17219)

x-x/3-x/4-x/6-x/10-x/12-x/20-1=0
x=60 oz not much of a crop
x/3=20 liberty
x/4=15 franklin
x/6=10 daffy
x/10=6 tony
x/12=5 jr
x/20=3 lady